# On the Optimality of Coded Distributed Computing for Ring Networks

### v1

@@ VI Conclusion
This paper investigates the tradeoff between NCL, computation load, and broadcast distance of the coded distributed systems over the ring network. For all-gather computing, we exactly characterized an information-theoretic tradeoff between NCL, computation load, and broadcast distance. For all-to-all computing, we proposed a coded transmission scheme that utilizes a simple network coding method for efficient broadcasting. We further prove that the proposed scheme is asymptotically optimal under cyclic data placement through information-theoretic converse arguments. We found that the coded transmission gain depends on the connectivity among nodes instead of redundant computation load in the network topology of the ring. Notably, the optimal tradeoff between NCL, computation load, and broadcast distance under arbitrary file placement remains an open problem for future work.
@@ Appendix A A Lower Bound on NCL for all-gather Computing
Appendix B A Lower Bound on NCL for all-to-all Computing under Cyclic Placement $(r\leq\lceil\frac{\mathsf{N}}{2}\rceil-1)$
Consider a cut as Fig.12 shown; we divide the nodes into two sets, $\mathcal{L}$ and the complement $\mathcal{R}$, where the nodes $\{n_{k-d},\ldots,n_{k}\}$ and $\{n_{j-d},\ldots,n_{j}\}$ belong to ${\cal L}$ have direct connection with ${\cal R}$. For any subset $\mathcal{S}\subseteq[\mathsf{N}]$ and a file assignment $\mathcal{M}$, we denote the set of IVs that are exclusively cached by $t$ nodes in $\cal{S}$ and desired by the nodes in $\cal{W}$ as $V_{t,\cal{S}}^{\cal{W}}$, and the number of these packets as $a_{t,\cal{S}}^{\cal{W}}$, i.e.,
$V_{t,\cal{S}}^{\cal{W}}\triangleq\big{(}\big{(}\underset{\begin{subarray}{c}\cal{T}\subseteq\cal{S}\\ |{\cal T}|={t}\\ k\in\cal{T}\end{subarray}}{\cap}{\cal M}_{k}\big{)}\backslash(\underset{i\notin{\cal S}}{\cup}{\cal M}_{k})\big{)}\cap\big{(}\underset{k\in{\cal W}}{\cap}{\cal W}_{k}\big{)},$ (34)
and $a_{t,\cal{S}}^{\cal{W}}\triangleq|V_{t,\cal{S}}^{\cal{W}}|$. Specifically, we denote the set IVs that are exclusively cached by nodes in ${\cal L}$ and desired by nodes in ${\cal R}$ as
$V_{\cal L}^{\cal R}=\{V_{t,\cal L}^{\cal R}:t\leq|{\cal L}|\}.$ (35)
Additionally, we denote the set of IVs that are cached by nodes in ${\cal R}$ as
=$\mathcal{M}_{\cal R}=\{v_{i}^{j}:i\in{\cal M}_{k},k\in{\cal R},j\in[\mathsf{N}]\},$ (36)
and denote $\{X_{k-d}^{\mathsf{T_{\text{e}}}},\ldots,X_{k}^{\mathsf{T_{\text{e}}}}\}$ and $\{X_{j-d}^{\mathsf{T_{\text{e}}}},\ldots,X_{j}^{\mathsf{T_{\text{e}}}}\}$ as $X_{kd}^{\mathsf{T_{\text{e}}}}$ and $X_{jd}^{\mathsf{T_{\text{e}}}}$, respectively.
Figure 12: A cut of the nodes. The nodes are divided into two complementary subsets, $\cal L$ and $\cal R$, separated by a dotted line.
@@
$H(V_{\cal L}^{\cal R})$ $=H(V_{\cal L}^{\cal R}|{\cal M}_{\cal R})$ (37)
$=I(X_{kd}^{\mathsf{T_{\text{e}}}},X_{jd}^{\mathsf{T_{\text{e}}}};V_{\cal L}^{\cal R}|{\cal M}_{\cal R})+H(V_{\cal L}^{\cal R}|X_{k}^{\mathsf{T_{\text{e}}}},X_{j}^{\mathsf{T_{\text{e}}}},{\cal M}_{\cal R})$ (38)
$=I(X_{kd}^{\mathsf{T_{\text{e}}}},X_{jd}^{\mathsf{T_{\text{e}}}};V_{\cal L}^{\cal R}|{\cal M}_{\cal R})+\mathsf{T_{\text{e}}}\epsilon_{\mathsf{T_{\text{e}}}}$ (39)
$=H(X_{kd}^{\mathsf{T_{\text{e}}}},X_{jd}^{\mathsf{T_{\text{e}}}}|{\cal M}_{\cal R})-H(X_{kd}^{\mathsf{T_{\text{e}}}},X_{jd}^{\mathsf{T_{\text{e}}}}|V_{\cal L}^{\cal R},{\cal M}_{\cal R})+\mathsf{T_{\text{e}}}\epsilon_{\mathsf{T_{\text{e}}}}$
$\leq H(X_{kd}^{\mathsf{T_{\text{e}}}},X_{jd}^{\mathsf{T_{\text{e}}}})+\mathsf{T_{\text{e}}}\epsilon_{\mathsf{T_{\text{e}}}}$ (40)
$=\sum_{i=1}^{d}\sum_{t=1}^{\mathsf{T_{\text{e}}}}(l_{k+1-i}^{(t)}+l_{j+1-i}^{(t)})+\mathsf{T_{\text{e}}}\epsilon_{\mathsf{T_{\text{e}}}},$ (41)
where (37) holds because of the independence between packets, (38) follows the definition of the mutual information, $\epsilon_{\mathsf{T_{\text{e}}}}$ is a function vanishing as $\mathsf{T_{\text{e}}}\to\infty$, (39) comes from that $H(V_{\cal L}^{\cal R}|Y_{\cal R}^{\mathsf{T_{\text{e}}}},{\cal M}_{\cal R})=\mathsf{T_{\text{e}}}\epsilon_{\mathsf{T_{\text{e}}}}$ holds because of Fano’s inequality and $Y_{\cal R}^{\mathsf{T_{\text{e}}}}$ are functions of $(X_{kd}^{\mathsf{T_{\text{e}}}},X_{jd}^{\mathsf{T_{\text{e}}}},{\cal M}_{\cal R})$. Besides,
$H(V_{\cal L}^{\cal R})=\sum_{t=1}^{|{\cal L}|}\sum_{k\in{\cal R}}a_{t,{\cal L}}^{\{k\}}\mathsf{B}_{2},$ (42)
@@
$\sum_{i=1}^{d}\sum_{t=1}^{\mathsf{T_{\text{e}}}}(l_{k+1-i}^{(t)}+l_{j+1-i}^{(t)})\geq\sum_{t=1}^{|{\cal L}|}\sum_{k\in{\cal R}}a_{t,{\cal L}}^{\{k\}}\mathsf{B}_{2}-\mathsf{T_{\text{e}}}\epsilon_{\mathsf{T_{\text{e}}}}.$ (43)
@@
Taking the partitions {$\{{\cal L},{\cal R}\}$ as the original version, which is referred to as $\{{\cal L}_{1},{\cal R}_{1}\}$, we can generate the new cut by moving some nodes. We move the node $n_{k+1}$ from ${\cal R}_{1}$ to ${\cal L}_{1}$, and the node $n_{j}$ from ${\cal L}_{1}$ to ${\cal R}_{1}$, leading to the new cut $\{{\cal L}_{2},{\cal R}_{2}\}$. There are $\mathsf{N}$ cuts like this, and the above arguments hold for every new cut. Summing up the (43) over all these cuts, we have
$2d\sum_{i=1}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}/\mathsf{B}_{2}\geq\sum_{i=1}^{\mathsf{N}}\sum_{t=1}^{|{\cal L}_{i}|}\sum_{k\in{\cal R}_{i}}a_{t,{\cal L}_{i}}^{\{k\}}-\epsilon^{\prime},$ (44)
@@
where $|{\cal L}_{1}|=\ldots=|{\cal L}_{\mathsf{N}}|$ and $\epsilon^{\prime}$ is a function vanishing as $\mathsf{T_{\text{e}}}\to\infty$. Note that (44) holds for $({\cal L}_{1},{\cal R}_{1},\ldots,{\cal L}_{\mathsf{N}},{\cal R}_{\mathsf{N}})$ of arbitrary size $|{\cal L}_{1}|=|{\cal L}|$. Recall that $a_{t,{\cal L}_{i}}^{\{k\}}$ denotes the number of the IVs that are exclusively cached by $t$ nodes in set ${\cal L}_{i}$ and intended for node $n_{k}$. Among these IVs, define the number of IVs that cached by nodes in set ${\cal G}$ of size $|{\cal G}|=t$ as $b_{{\cal G},{\cal L}_{i}}$, which implies that $\sum_{\begin{subarray}{c}{\cal G}\subseteq{\cal L}_{i}\\ |{\cal G}|=t\end{subarray}}b_{{\cal G},{\cal L}_{i}}=a_{t,{\cal L}_{i}}^{\{k\}}$. We have
$\sum_{i}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}/\mathsf{B}_{2}+\frac{\epsilon^{\prime}}{2d}$ $\geq\frac{1}{2d}\cdot\max_{|{\cal L}|}\left\{\sum_{i}^{\mathsf{N}}\sum_{t=1}^{|{\cal L}|}\sum_{k\in{\cal R}_{i}}a_{t,{\cal L}_{i}}^{\{k\}}\right\}$
$\geq\frac{1}{2d}\sum_{i}^{\mathsf{N}}\sum_{t=1}^{|{\cal L}|=\frac{\mathsf{N}}{2}}|{\cal R}_{i}|a_{t,{\cal L}_{i}}^{\{k\}}$ (45)
$=\frac{\mathsf{N}}{4d}\sum_{i}^{\mathsf{N}}\sum_{t=1}^{|{\cal L}|=\frac{\mathsf{N}}{2}}a_{t,{\cal L}_{i}}^{\{k\}}$ (46)
$=\frac{\mathsf{N}}{4d}\sum_{i}^{\mathsf{N}}\sum_{t=1}^{|{\cal L}|=\frac{\mathsf{N}}{2}}\sum_{\begin{subarray}{c}{\cal G}\subseteq{\cal L}_{i}\\ |{\cal G}|=t\end{subarray}}b_{{\cal G},{\cal L}_{i}}$ (47)
For the cyclic placement,
$\sum_{t=1}^{|{\cal L}|=\frac{\mathsf{N}}{2}}\sum_{\begin{subarray}{c}{\cal G}\subseteq{\cal L}_{i}\\ |{\cal G}|=t\end{subarray}}b_{{\cal G},{\cal L}_{i}}=\sum_{\begin{subarray}{c}{\cal G}\subseteq{\cal L}_{i}\\ |{\cal G}|=r\end{subarray}}b_{{\cal G},{\cal L}_{i}}=\frac{\mathsf{N}}{2}-r+1$ (48)
when $r<\frac{\mathsf{N}}{2}$, which implies that
$\sum_{i}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}/(\mathsf{N}\mathsf{B}_{2})$ $\geq\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}-r+1)-\epsilon^{\prime}$
$=\frac{(\mathsf{N}-r)^{2}}{8d}+\frac{\mathsf{N}-\frac{r^{2}}{2}}{4d}-\epsilon^{\prime}.$ (49)
@@ Appendix C The optimal NCL for All-to-all When $r\geq\frac{\mathsf{N}}{2}$ and $d=1$
achieve the NCL of $\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil.$ The proposed transmission strategy for the problem is straightforward. We provide an example to illustrate it. The file placement for $\mathsf{N}=8$, $r=4$ is shown as Table V. The symbol * at the row $n_{i}$ and column $w_{j}$ where $i\in[8]$ and $j\in[8]$ means that the node $n_{i}$ compute the IVs from $w_{j}$. Node $n_{i}$ broadcast $X_{i}^{(1)}=v_{i}^{i+1}\oplus v_{i+1}^{i-1}$ and $X_{i}^{(2)}=v_{i+4}^{i+1}\oplus v_{i+5}^{i-1}$. For example, node $n_{4}$ broadcasts
@@
TABLE V: A file placement achieving optimal transmission for $\mathsf{N}=8$, $r=4$ and $d=1$
@@
where $\epsilon_{\prime}$ vanishes as $\mathsf{T_{\text{e}}}\to\infty$. This completes the proof.

### v2

@@ VI Concluding Remark and Future Directions
This paper investigates the tradeoff between NCL, computation load, and broadcast distance of the coded distributed systems over the ring network. For all-gather computing, we exactly characterized an information-theoretic tradeoff between NCL, computation load, and broadcast distance. For all-to-all computing, we proposed a coded transmission scheme that utilizes a simple network coding method for efficient broadcasting. We further prove that the proposed scheme is asymptotically optimal under cyclic data placement through information-theoretic converse arguments. We found that the coded transmission gain depends on the connectivity among nodes instead of redundant computation load in the network topology of the ring. Notably, the optimal tradeoff between NCL, computation load, and broadcast distance under arbitrary file placement remains an open problem for future work. Besides, extensions to other network topologies as follows, such as 2-D torus networks, satellite networks, are also worth investigating.
VI-A Extension to $2$-d Torus Networks
The ring network studied in this work is a one-dimensional $k$-nearest-neighbor circulant graph. The proposed framework can be extended to other structured topologies, such as two-dimensional torus, provided that the broadcast capability is defined in a compatible manner. In particular, consider a 2-D torus network where each node performs distance-limited broadcast along one dimension at a time, with the broadcast distance defined analogously to the ring model. In this case, a natural extension is a dimension-wise nested scheme. For the all-gather task, one may first apply the proposed ring-based coded transmission independently along one dimension (e.g., rows) to aggregate intermediate values within each row, and then apply the same scheme along the other dimension (e.g., columns) to complete global aggregation. For the all-to-all task, a similar nesting applies, where intermediate values are first delivered to their target rows and then delivered to their final destinations along the second dimension using the proposed schedules within each sub-ring. Under this nested construction, reverse-carpooling opportunities arise independently along each dimension, preserving the key insights established for ring networks. If the broadcast model differs substantially from that assumed above, new coding and scheduling designs may be required, which is a worthy direction for future work.
VI-B Relevance to Satellite Orbital-Plane Networks.
The ring abstraction studied in this work is also relevant to satellite constellations, where satellites within a single orbital plane form a ring via inter-satellite links (ISLs) [23, 24]. In this setting, $\mathsf{N}$ corresponds to the number of satellites per plane, $d$ represents the hop budget of ISL communication within a scheduling window, and $r$ reflects redundancy enabled by replicated sensing or preprocessing. The all-gather and all-to-all tasks naturally arise from plane-wide aggregation (e.g., ephemeris dissemination or onboard model aggregation) and directional exchanges (e.g., task offloading or cooperative calibration), respectively. The reduction in normalized communication load implies fewer ISL transmissions, which is particularly relevant under half-duplex and slotted operation. The investigations under more complex broadcast or dynamic connectivity models are interesting future directions.
@@ Appendix A A Lower Bound on NCL for all-gather Computing
Appendix B A Lower Bound on NCL for all-to-all Computing $(r\leq\lceil\frac{\mathsf{N}}{2}\rceil-1)$
Consider a cut as Fig. 13 shown; we divide the nodes into two sets, $\mathcal{A}_{j}$ and the complement $\bar{\mathcal{A}}_{j}$, where the nodes $\{n_{j-d},\ldots,n_{j-1}\}$ and $\{n_{j+s},\ldots,n_{j+s+d-1}\}$ belong to $\bar{\cal A}_{j}$} have direct connection with ${\cal A}_{j}$. Define the set of files cached by nodes in ${\cal A}_{j}$ as
$\mathcal{M}_{{\cal A}_{j}}\triangleq\underset{{k\in{\cal A}_{j}}}{\cup}{\cal M}_{k}\subseteq\{w_{1},\ldots,w_{\mathsf{N}}\},$ (39)
and its complement as
$\bar{\mathcal{M}}_{{\cal A}_{j}}\triangleq\{w_{1},\ldots,w_{\mathsf{N}}\}\backslash\mathcal{M}_{{\cal A}_{j}}$ (40)
which has the size of $|\bar{\mathcal{M}}_{{\cal A}_{j}}|=m_{j}$. Additionally, we denote the set of IVs that are required by nodes in ${\cal A}_{j}$ but cannot be computed locally inside ${\cal A}_{j}$ as
$V({\cal A}_{j})=\{v_{i}^{k}:w_{i}\in\bar{\mathcal{M}}_{{\cal A}_{j}},k\in{\cal A}_{j}\},$ (41)
and denote $\{X_{j-d}^{\mathsf{T_{\text{e}}}},\ldots,X_{j-1}^{\mathsf{T_{\text{e}}}}\}$ and $\{X_{j+s}^{\mathsf{T_{\text{e}}}},\ldots,X_{j+s+d-1}^{\mathsf{T_{\text{e}}}}\}$ as $X_{j-}^{\mathsf{T_{\text{e}}}}$ and $X_{j+}^{\mathsf{T_{\text{e}}}}$, respectively.
[…]
@@ Appendix B A Lower Bound on NCL for all-to-all Computing $(r\leq\lceil\frac{\mathsf{N}}{2}\rceil-1)$
$H(V({\cal A}_{j}))$ $=H(V({\cal A}_{j})|\mathcal{M}_{{\cal A}_{j}})$ (42)
$=I(X_{j-}^{\mathsf{T_{\text{e}}}},X_{j+}^{\mathsf{T_{\text{e}}}};V({\cal A}_{j})|\mathcal{M}_{{\cal A}_{j}})+H(V({\cal A}_{j})|X_{j-}^{\mathsf{T_{\text{e}}}},X_{j+}^{\mathsf{T_{\text{e}}}},\mathcal{M}_{{\cal A}_{j}})$ (43)
$=I(X_{j-}^{\mathsf{T_{\text{e}}}},X_{j+}^{\mathsf{T_{\text{e}}}};V({\cal A}_{j})|\mathcal{M}_{{\cal A}_{j}})$ (44)
$=H(X_{j-}^{\mathsf{T_{\text{e}}}},X_{j+}^{\mathsf{T_{\text{e}}}}|\mathcal{M}_{{\cal A}_{j}})-H(X_{j-}^{\mathsf{T_{\text{e}}}},X_{j+}^{\mathsf{T_{\text{e}}}}|V({\cal A}_{j}),\mathcal{M}_{{\cal A}_{j}})$
$\leq H(X_{j-}^{\mathsf{T_{\text{e}}}},X_{j+}^{\mathsf{T_{\text{e}}}})$ (45)
$=\sum_{i=1}^{d}\sum_{t=1}^{\mathsf{T_{\text{e}}}}(l_{j-i}^{(t)}+l_{j+s+i-1}^{(t)}),$ (46)
where (42) holds because of the independence between packets, (43) follows the definition of the mutual information, (44) comes from that $H(V({\cal A}_{j})|Y_{{\cal A}_{j}}^{\mathsf{T_{\text{e}}}},\mathcal{M}_{{\cal A}_{j}})=0$ holds because of the decoding constraint and $Y_{{\cal A}_{j}}^{\mathsf{T_{\text{e}}}}$ are functions of $(X_{j-}^{\mathsf{T_{\text{e}}}},X_{j+}^{\mathsf{T_{\text{e}}}},\mathcal{M}_{{\cal A}_{j}})$, and (45) holds because conditioning reduces entropy, and (46) follows the definition of $X_{j-}^{\mathsf{T_{\text{e}}}}$ and $X_{j+}^{\mathsf{T_{\text{e}}}}$. Besides,
$H(V({\cal A}_{j}))=|\bar{\mathcal{M}}_{{\cal A}_{j}}|\cdot|{\cal A}_{j}|\cdot\mathsf{B}_{2}=m_{j}\cdot s\cdot\mathsf{B}_{2},$ (47)
@@
$\sum_{i=1}^{d}\sum_{t=1}^{\mathsf{T_{\text{e}}}}(l_{j-i}^{(t)}+l_{j+s+i-1}^{(t)})\geq m_{j}\cdot s\cdot\mathsf{B}_{2}.$ (48)
@@
Taking the partitions $\{{\cal A}_{j},\bar{\cal A}_{j}\}$ as the original version, we can generate the new cut by moving some nodes. We move the node $n_{j}$ from ${\cal A}_{j}$ to $\bar{\cal A}_{j}$}, and the node $n_{j+s}$ from $\bar{\cal A}_{1}$} to ${\cal A}_{1}$, leading to the new cut $\{{\cal A}_{j+1},\bar{\cal A}_{j+1}\}$. There are $\mathsf{N}$ cuts like this, and the above arguments hold for every new cut. Summing up the above inequality (48) over all these cuts, we have
$\sum_{j=1}^{\mathsf{N}}\sum_{i=1}^{d}\sum_{t=1}^{\mathsf{T_{\text{e}}}}(l_{j-i}^{(t)}+l_{j+s+i-1}^{(t)})\geq\sum_{j=1}^{\mathsf{N}}m_{j}\cdot s\cdot\mathsf{B}_{2},$ (49)
which is equivalent to
$2d\sum_{i=1}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}\geq\sum_{j=1}^{\mathsf{N}}m_{j}\cdot s\cdot\mathsf{B}_{2}.$ (50)
@@
Note that (49) holds for $({\cal A}_{1},\bar{\cal A}_{1},\ldots,{\cal A}_{\mathsf{N}},\bar{\cal A}_{\mathsf{N}})$ of arbitrary size $|{\cal A}_{1}|=\ldots=|{\cal A}_{\mathsf{N}}|=s$. Recall that $m_{j}$ denotes the number of files that are not cached by nodes in ${\cal A}_{j}$. We have
$\sum_{j=1}^{\mathsf{N}}|{\cal{M}}_{{\cal A}_{j}}|\leq\sum_{j=1}^{\mathsf{N}}\sum_{u\in{\cal A}_{j}}|{\cal M}_{u}|\leq s\sum_{u=1}^{\mathsf{N}}|{\cal M}_{u}|=sr\mathsf{N},$ (51)
which implies that
$\sum_{j=1}^{\mathsf{N}}m_{j}=\sum_{j=1}^{\mathsf{N}}(\mathsf{N}-|{\cal{M}}_{{\cal A}_{j}}|)\geq\mathsf{N}^{2}-sr\mathsf{N}.$ (52)
Then, we have
$\sum_{i=1}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}/(\mathsf{N}\mathsf{B}_{2})$ $\geq\max_{s\in\{1,\ldots\mathsf{N}\}}\frac{\sum_{j=1}^{\mathsf{N}}m_{j}\cdot s\cdot\mathsf{B}_{2}}{2d\cdot\mathsf{N}\mathsf{B}_{2}}$
$\geq\max_{s\in\{1,\ldots\mathsf{N}\}}\frac{s(\mathsf{N}-sr)}{2d}.$ (53)
  • For the arbitrary file placement given computation load $r$, the optimizer is typically around $s=\frac{\mathsf{N}}{2r}$ (requiring $s\geq 1$, i.e, $r\leq\frac{\mathsf{N}}{2}$), giving the scaling of lower bound as
    $\sum_{i=1}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}/(\mathsf{N}\mathsf{B}_{2})\gtrsim\frac{\mathsf{N}^{2}}{8dr}.$ (54)
  • For the cyclic file placement given computation load $r$, we have $m_{j}=\mathsf{N}-s-r_{j}+1$ for all $j\in[\mathsf{N}]$, which refines the (52) as
    $\sum_{j=1}^{\mathsf{N}}m_{j}=\sum_{i=1}^{\mathsf{N}}(\mathsf{N}-s-r_{i}+1)=\mathsf{N}^{2}-s\mathsf{N}-r\mathsf{N}+\mathsf{N},$ (55)
    where $r_{i}$ is the number of nodes that cache file $w_{i}$. Then, we have
    $\sum_{i=1}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}/(\mathsf{N}\mathsf{B}_{2})\geq\max_{s\in\{1,\ldots\mathsf{N}\}}\frac{s(\mathsf{N}-s-r+1)}{2d},$ (56)
    which is optimized at $s=\frac{\mathsf{N}-r+1}{2}$ (requiring $s\geq 1$, i.e, $r\leq\mathsf{N}-1$), giving the scaling of lower bound as
    $\sum_{i=1}^{\mathsf{N}}\sum_{t=1}^{\mathsf{T_{\text{e}}}}l_{i}^{(t)}/(\mathsf{N}\mathsf{B}_{2})\gtrsim\frac{(\mathsf{N}-r+1)^{2}}{8d}.$ (57)
@@ Appendix C The optimal NCL for All-to-all When $r\geq\frac{\mathsf{N}}{2}$ and $d=1$
achieve the NCL of $\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil.$ The proposed transmission strategy for the problem is straightforward. We provide an example to illustrate it. The file placement for $\mathsf{N}=8$, $r=4$ is shown as Table VI. The symbol * at the row $n_{i}$ and column $w_{j}$ where $i\in[8]$ and $j\in[8]$ means that the node $n_{i}$ compute the IVs from $w_{j}$. Node $n_{i}$ broadcast $X_{i}^{(1)}=v_{i}^{i+1}\oplus v_{i+1}^{i-1}$ and $X_{i}^{(2)}=v_{i+4}^{i+1}\oplus v_{i+5}^{i-1}$. For example, node $n_{4}$ broadcasts
@@
TABLE VI: A file placement achieving optimal transmission for $\mathsf{N}=8$, $r=4$ and $d=1$
@@
where $\epsilon^{\prime}$ vanishes as $\mathsf{T_{\text{e}}}\to\infty$. This completes the proof.